OpenStudy (anonymous):

What is the solution of the following system? -3x-2y=-12 and 9x+6y=-9

OpenStudy (texaschic101):

-3x - 2y = -12 -->(3) 9x + 6y = -9 --------------- -9x - 6y = -36 (result of multiplying by 3) 9x + 6y = -9 ---------------add 0 = - 45 (incorrect) This problem has NO SOLUTION